OpenStudy (anonymous):

Which lines are perpendicular to 7x – 3y = 18? Choose all answers that are correct. (Points : 1) 3x + 7y = 42 3y = 7x – 15 3x – 7y = 21 –7y = 3x + 1

OpenStudy (godlovesme):

first change all equations to slope-intercept formula @FlowerChald

OpenStudy (anonymous):

ok ill try to do that

OpenStudy (anonymous):

the formula is y=mx+b right?

OpenStudy (godlovesme):

yes

OpenStudy (anonymous):

how excatly would i change the answers into intercept form?

OpenStudy (igreen):

Just rearrange the equations. Slope intercept form is y = mx + b.

OpenStudy (godlovesme):

u wanna get y alone so just solve for y

OpenStudy (igreen):

3x + 7y = 42 Subtract 3x to both sides: 7y = -3x + 42 Divide 7 to both sides: y = -3/7x + 6 Okay, now you try the rest.

OpenStudy (anonymous):

ok thank you